OpenStudy (anonymous):

find the sum of series 3^(1/2)+3+3*3^(1/2)+.... n terms

OpenStudy (anonymous):

\[1+\sqrt{3}+3+3\sqrt{3}+9+9\sqrt{3}+....... infinite\]

hartnn (hartnn):

do you realize both of these are geometric series ? if yes, can you tell me common ratio for both ?

OpenStudy (anonymous):

yup i realise it but as much as i remember sum of infinite series is 1/(1-r) where \[\left| r \right|\le1\]

OpenStudy (anonymous):

here r=\[\sqrt{3}\]

OpenStudy (anonymous):

and \[\sqrt{3}>1\]

hartnn (hartnn):

yes, so that means the sum of this series is INFINITE too!

OpenStudy (anonymous):

but do you notice that r>1 which means formula is not applicable

hartnn (hartnn):

yes, when r>1 , sum is infinite when number of terms are infinite simply because sum goes on incrasing...

OpenStudy (anonymous):

ok thats why the answer was not matching thanks by the way so it means answer is simply infinte

hartnn (hartnn):

yes!